40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 5322     Entry time: Tue Aug 30 10:49:29 2011
Author: steve 
Type: Update 
Category: SUS 
Subject: BS & PRM damping restored 

I have restored the damping of BS and PRM. Today is janitor day. He is shaking things around the lab.

ELOG V3.1.3-